Butler Corporation issued 15-year bonds three years ago at a coupon rate of 8% and a face value of $1000. The bonds make semiannual payments. If these bonds currently sell for $1050 what is the annual yield to maturity? please show work
Answer #1

Coupon amount = 1000*8%/2 = 40 PV = 1050 FV = 1000 Time to maturity = 15-3 = 12*2 = 24 Using excel rate function: Yield to ma